The measure of the larger acute angle in a right triangle is two degrees less than three times the measure of the smaller acute angle. Find the measure of each angle. Please help me!!! atleast find the equation:) thanx
Answer by checkley77(12844)

(3x-2)+x=90
3x-2+x=90
4x=90+2
4x=92
x=92/4
x=23 degrees is the smaller angle.
90-23=67 is the larger angle.
Proof:
(3*23-2)+23=90
(69-2)+23=90
67+23=90
90=90